A jar contains 8 pennies, 4 nickels and 4 dimes. A child selects 2 coins at random without replacement from the jar. Let X represent the amount in cents of the selected coins. Find the probability X=10


1
Expert's answer
2021-02-24T13:23:43-0500

Given coins are , 8 pennies, 4 nickels, 4 dimes


As X denote the amount of cents in selected coins

"10 \text{ cents}=2 \text{ nickels}"


So probability of (X=10), contains withdrwal of 2 nickels without replacement

Probability of drawing first nickel ="\dfrac{4}{16}"


Probability of drawing second nickel ="\dfrac{3}{15}"


"\implies P(X=10)=(\dfrac{4}{16})(\dfrac{3}{15})=\dfrac{1}{20}"


Hence, Probability "(X=10)" is "\dfrac{1}{20}."